OpenStudy (anonymous):

Gino designed a garden for his rectangular backyard. The design he drew on paper is ten inches long. The length of his backyard is fifty feet. If his drawing is to scale and is 9 inches wide, how many feet wide is his backyard? Write a proportion using words. Then, use the proportion to calculate the yard width. Show your work.

OpenStudy (anonymous):

Alright Mumina, after reading this, what are you trying to solve for?

OpenStudy (campbell_st):

convert your meaurements to the same units 50 ft = 600 inches so 10:600 or 1:60 then \[\frac{600}{10} = \frac{x}{9}\] then x = 540 inches x = 540/12 the width of the garden is 45 feet